Fred jogs uphill at 8mph to work.On his way home (downhill) he averages 11mph and takes him 15 min less time.How far does he live from work?
Answer by checkley77(12844)

8t=11(t-.25)
8t=11t-2.75
8t-11t=-2.75
-3t=-2.75
t=-2.75/-3
t=.917 hours is his time.
8*.917=7.33 miles is the distance to his work.
proof
11(.917-.25)=11*.667=7.33 mles.